OpenStudy (mckenzieandjesus):

Find the first, fourth, and eighth terms of the sequence. A(n)=-5*2^n^-1 –5; –40; –640 –20; –160; –2,560 –10; –80; –1,280 1; –1,000; –10,000,000

OpenStudy (mckenzieandjesus):

@Mehek14 can please u help?

OpenStudy (mckenzieandjesus):

@Loser66 can u help?

OpenStudy (mckenzieandjesus):

@jim_thompson5910

OpenStudy (mckenzieandjesus):

@hartnn can u please help?

OpenStudy (marigirl):

\[A(n)=5\times 2^{n-1}\] Is that how the equation is showing?

OpenStudy (mckenzieandjesus):

yes!

OpenStudy (marigirl):

So first term you would sub in n=1 \[A(n)= -5 \times 2^{n-1}\] \[A(1)= -5 \times 2^{1-1}\] \[A(1)= -5 \times 2^{0}\] \[A= -5 \times 1\] A=-5

OpenStudy (marigirl):

fourth sub in n=4 \[A= -5 \times 2^{4-1}\] \[A= -5 \times 2^{3}\] \[A= -5 \times 8\] A=-40

OpenStudy (marigirl):

try the eight term and see if you find the matching sequence from the answers :)

OpenStudy (mckenzieandjesus):

-640

OpenStudy (mckenzieandjesus):



OpenStudy (mckenzieandjesus):

Thank You so much. Could you help me with this one? Evaluate the function rule for the given value. f(x)=15^x for x=3

OpenStudy (mckenzieandjesus):

15^3=3375

OpenStudy (marigirl):

yes that looks good